OpenStudy (anonymous):

this year the high school basketball team boasted a winning season by winning one omore game than it lost. the team played a total of 23 games. how many games did the team win?

OpenStudy (ganpat):

u will get 2 equations.. x = games won and y = games lost 1. x + y = 23 2. x = y + 1 solve them, and you will have answer

OpenStudy (ganpat):

y + 1 + y = 23 2y = 22 y = 11 and x = 12

OpenStudy (anonymous):

thank you